if A(Z 1),B(Z2),C(Z3) on complex plane such that angles at B and C are each equal to 1/2(∏-α) then prove that (Z2-Z3) 2 =4(Z3 - Z1)(Z1-Z2)sin 2 α/2

if A(Z1),B(Z2),C(Z3) on complex plane such that angles at B and C are each equal to 1/2(∏-α) then prove that (Z2-Z3)2 =4(Z3 - Z1)(Z1-Z2)sin2 α/2

Grade:11

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear deepika


<B =<c =(∏-α)/2    so angel A = α

now use sine rule in trinagle ABC

|z2 -z3|/sinα  = |z1-z3|/cosα = |z1-z2|/cosα

now by the rotation property of vector

(z2-z3)/|z2 -z3|  = (z1-z3)/|z1-z3|  ei(∏-α)/2   ...........1


and  also (z2-z3)/|z2 -z3| = (z2-z1)/|z2-z1|  ei(∏-α)/2   ...........2


multiple first and second equation

(Z2-Z3)2  /|z2 -z3|2=(Z3 - Z1)(Z1-Z2)/|z1-z3||z2-z1|


now put the value of |z1-z3|, |z2-z1| in terms of |z2 -z3|

(Z2-Z3)2  =sin2α(Z3 - Z1)(Z1-Z2)/cos2α/2

(Z2-Z3)2 =4sin2α/2 (Z3 - Z1)(Z1-Z2)

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin